A boat travels down the river for 8 miles in 20 minutes. The same boat travels
back up the river for 8 miles in 30 minutes. Find the speed of the boat in
still water.

Answer by stanbon(75887)

A boat travels down the river for 8 miles in 20 minutes.
The same boat travels back up the river for 8 miles in 30 minutes.
Find the speed of the boat in still water.
-----------------------------------------------
Downriver DATA:
distance = 8 miles ; time = (1/3)hr. ; rate = d/t = 8/(1/3) = 24 mph
----------------------------------------------
Upriver Data:
distance = 8 miles; time = (1/2) hr ; rate = d/t = 8/(1/2) = 16 mph
-----------------------------------------------
Equations:
b + c = 24
b - c = 16
-------------
2b = 40
b = 20 mph (speed of the boat in still water)
----
Substitute to solve for "c":
20 + c = 24
c = 4 mph (speed of the current)
=======================================
Cheers,
Stan H.